OpenStudy (anonymous):

Admission to a baseball game game is $3.00 for general admission and $3.50 for reserved seats. The receipts were $2907.50 for 906 admissions. How many of each ticket were sold? (round to the nearest integer if necessary)

OpenStudy (anonymous):

We need to create s system of equations to solve. For the first equation, lets look at the number of tickets sold. Let x represent the general seats, and y represent the reserved seats. Then we know that: x + y = 906 because there was 906 addmissions. Now, looking at the price instead, we have: 3.00x + 3.50y = 2907.50 because 2907.50 was spent in total on those tickets.

OpenStudy (anonymous):

So now, we need to solve this system, and there are a couple of ways to do that. Im going to do it by substitution, but you can do it in whatever method you please. So looking at the first equation, im going to solve for x: \[x+y = 906 \Rightarrow x = 906-y\] Now im going to substitute that value of x into the second equation, and solve for y: \[3.00x+3.50y=2907.50 \Rightarrow 3.00(906-y)+3.50y = 2907.50\] \[\Rightarrow 2718.00-3.00y+3.50y = 2907.50 \Rightarrow .50y = 189.50 \Rightarrow y = 379\]

OpenStudy (anonymous):

Now, y represented the number of reserved seats, so we still need to figure out the number of general admission seats. That would be: 906 - 379 = 527 and now you are done :)